OpenStudy (anonymous):

∫ (secxtanx)/(4+sec^2x) dx Please show steps on what to do, consider inverse trigonometric functions. Should I make a u-substitution?

hartnn (hartnn):

if you put u= sec x , what will be du= ... ?

OpenStudy (anonymous):

secxtanx would be the du?

hartnn (hartnn):

yes, that correct. so substitute these in your integral, what will your new integral in terms of u and du ?

OpenStudy (anonymous):

Give me a moment to write this out

hartnn (hartnn):

and actually, du = sec x tan x dx

hartnn (hartnn):

sure, take your time :)

OpenStudy (anonymous):

Wouldn't having the u=4+sec^2x be easier? and then it would just be du=2secxtanx?

OpenStudy (anonymous):

That way it'll cancel with the first secxtanx

hartnn (hartnn):

u= 4+sec^2x du = 2 sec x (sec x tan x) dx not easier.

OpenStudy (anonymous):

So my dx= 1/2 du secxtanx?

OpenStudy (anonymous):

Alright so I have ∫ (secxtanx) (4)+ (u)^2 dx where u=secx? Or does that first secxtanx get cancelled out somehow..

hartnn (hartnn):

lol, don't confuse me \(\large \int \dfrac{(\sec x \tan xdx)}{4+\sec^2x}=\int \dfrac{du}{4+u^2}\) because numerator = sec x tan x dx = du

OpenStudy (anonymous):

Lol I had the 4+sec^2x up top so it would be (4+sec^2x)^-1

OpenStudy (anonymous):

Don't I have to solve for dx?

hartnn (hartnn):

means ? do what you think needs to be done, and i'll check/verify. also, didn't you get the step i did ?

OpenStudy (anonymous):

Ok so I have the∫ du/(4+(u)^2) Would I just anti-derive it to be 1/4 * ln(u)^2?

hartnn (hartnn):

no... antiderivative of 1/(x^2+a^2) is 1/a tan^{-1} (x/a) +c heard of it ?

OpenStudy (anonymous):

No, no I havent.

hartnn (hartnn):

now you have :)

OpenStudy (anonymous):

I'm still so very lost... :/ I'm stuck at ∫ du/ 4+u^2 x still

OpenStudy (anonymous):

Where is this x^2 coming from in relation to my problem?

hartnn (hartnn):

doesn't these 2 look similar ? \(\huge \int \dfrac{1}{x^2+a^2}dx= \dfrac{1}{2}\tan^{-1}(\dfrac{x}{a})+c\) and \(\huge \int \dfrac{1}{u^2+4}du=....?\) you must be knowing, 4= ...^2

OpenStudy (anonymous):

No I never learned that. My teacher never showed us that formula.

OpenStudy (anonymous):

So the final answer would be 1/2 (tan^-1) (sec^2)/(4) +C?

hartnn (hartnn):

a= 4 or 2 ?

hartnn (hartnn):

and u = sec x only.

OpenStudy (anonymous):

2

hartnn (hartnn):

and sorry, the formula is \(\huge \int \dfrac{1}{x^2+a^2}dx= \dfrac{1}{a}\tan^{-1}(\dfrac{x}{a})+c\)

OpenStudy (anonymous):

I'm just trying to work this out in a way my teacher would deem fit since we haven't learned that formula

hartnn (hartnn):

ok, so lets prove it using a trigonometric substitution u =2tan y du = sec^2 y dy \(\huge \int \dfrac{1}{u^2+4}du=\int \dfrac{\sec^2 ydy}{4\tan^2 y+4}\) can you simplify the denominator ??

OpenStudy (anonymous):

There is no 2tanx in the problem though so how can that be u?

hartnn (hartnn):

hmmm.. ? i am making a new substitution in 1/ (u^2+4) du u= 2 tan y didn't get what you doubt is ?

OpenStudy (anonymous):

Alright let's back track here and etch a sketch the thing. Tell me if this is just totally wrong b/c this is the way my brain wants to work. ∫ (secxtanx)(4+sec^2x)^-1 u=4+sec^2x du=2secx+secxtanx dx Integrating it ∫ (2secx)/(u) (2secx)/(4+sec^2x) final answer. That is how my teacher taught us but I feel like I'm going wrong somewhere in there or what.

OpenStudy (anonymous):

The first secxtanx cancel out b/c of the secxtanx in the du

hartnn (hartnn):

du=2secx+secxtanx dx <----incorrect.

OpenStudy (anonymous):

What is the du then, just secxtanx? I though I had to use the chain rule for that.

hartnn (hartnn):

du = 2 sec x (sec x tan x ) dx by chain rule. this du is of no use here.

hartnn (hartnn):

if u= sec x then du = sec x tan x dx which is actually the numerator!

OpenStudy (anonymous):

My poor eraser, okay let me figure it out that way.

zepdrix (zepdrix):

\[\large \color{orangered}{u=\sec x}\]\[\large \color{royalblue}{du=\sec x \tan x dx}\] \[\large \int\limits\limits \frac{\color{royalblue}{\sec x \tan x dx}}{4+(\color{orangered}{\sec x})^2} \qquad \rightarrow \qquad \int\limits \frac{\color{royalblue}{du}}{4+(\color{orangered}{u})^2}\] Maybe colors will help sort things out. I always find that helps me at least :D

OpenStudy (anonymous):

∫ du/4+(u)2 dx okay, and how do I find the anti derivative of that or is that just it?

zepdrix (zepdrix):

You have two options, you can either: ~Make a trigonometric substitution Or ~Just memorize the formula, depending on which method you have learned at this point.

OpenStudy (anonymous):

Yes that's where I'm stuck, the second part there.

OpenStudy (anonymous):

do the trigonometric substitution please

zepdrix (zepdrix):

The point of making a trig sub is that it allows us to take advantage of our trig identities. We can simplify the denominator to one term, getting rid of the addition. Since it's of the form \(\large a^2+x^2\) we'll make the substitution \(\large x=a\tan\theta\) Here is what that will do for us. \[\large a^2+x^2 \qquad \rightarrow \qquad a^2+(a \tan \theta)^2 \qquad \rightarrow \qquad a^2(1+\tan^2\theta)\] From here, we revisit our Trig Identity, \(\large \color{royalblue}{1+\tan^2\theta=\sec^2\theta}\) Which gives us, \(\large a^2\sec^2\theta\) See how that gets rid of the addition? This term that it leaves us with is much easier to integrate. So let's rewrite our integral a sec, so we have an \(\large a^2\) in the bottom.

OpenStudy (anonymous):

I never learned that formula though

OpenStudy (anonymous):

the a^2+x^2 thing

zepdrix (zepdrix):

You didn't? Then why are you working on this problem? Have you learned this formula then? \[\large \int\limits \frac{1}{a^2+x^2}dx \qquad = \qquad \frac{1}{a}\arctan\left(\dfrac{x}{a}\right)\]

OpenStudy (anonymous):

I have not -_- I'm just going to say screw it and learn it anyhow. Apparently my teacher left something out or I missed something. I'll just go with it. Okay, so re-writing the problem with the a^2(sec^2θ).. continue

OpenStudy (anonymous):

1/2 (arctan)*sec/2 ?

zepdrix (zepdrix):

We first want to get the proper form in the denominator.\[\large \int\limits \frac{1}{4+u^2}du \qquad = \qquad \int\limits \frac{1}{2^2+u^2}du\] Ok now we can properly see our \(\large a\) value is \(\large 2\). We'll make the substitution,\[\large u=a \tan \theta \qquad \rightarrow \qquad \color{royalblue}{u=2\tan \theta}\]As with any substitution, we need to take a derivative, so we can replace the differential \(\large du\). \(\large \color{orangered}{du=2\sec^2\theta}\) Plugging in all the pieces,\[\large \int\limits\limits \frac{\color{orangered}{du}}{2^2+\color{royalblue}{u}^2} \qquad =\qquad \int\limits\limits \frac{\color{orangered}{\color{orangered}{2\sec^2\theta}}}{2^2+(\color{royalblue}{2\tan \theta})^2} \]

zepdrix (zepdrix):

That last orange term should have a \(\large d\theta\), my bad. So this will simplify to,\[\large \frac{2}{4}\int\limits \frac{\sec^2\theta d \theta}{1+\tan^2\theta}\] Applying our Trig Identity to the denominator gives us,\[\large \frac{1}{2}\int\limits\limits \frac{\sec^2\theta\; d \theta}{\sec^2\theta} \qquad =\qquad \frac{1}{2}\int\limits d \theta\]

zepdrix (zepdrix):

Trig Substitution can be a little tricky to get used to. With U-substitutions, you're replacing a big chunk of stuff with a single variable. With Trig Substitutions, you're replacing something with a `function` of a single variable. The function being the Trig function. So it might seem a bit strange at first.

zepdrix (zepdrix):

FINALLY integrating gives us,\[\large \frac{1}{2}\int\limits d \theta \qquad = \qquad \frac{1}{2}\theta+C\] The final steps will be rewriting our answer in terms of \(\large x\). To do so, we have to back-substitute into \(\large u\). From there, back-substitute again into \(\large x\).

OpenStudy (anonymous):

Alright, I had someone else solve the problem to equal (1/2)arctan(sec(x)/2) + C. Does this seem correct or no?

zepdrix (zepdrix):

Yah that looks right :) Too confusing I guess? Heh.

OpenStudy (anonymous):

Oh boy. Now I have no idea where to go lol

zepdrix (zepdrix):

Wanna see the last steps or not so much? :O Too much to take in?

OpenStudy (anonymous):

I'm not sure which one is the correct answer now..

OpenStudy (anonymous):

What do you think?

zepdrix (zepdrix):

I don't understand what you're asking. The answer you posted is correct. I hadn't finish the problem yet... Where are you seeing conflicting answers...?

OpenStudy (anonymous):

Where did dθ come from in yours

OpenStudy (anonymous):

Is that in place of du?

zepdrix (zepdrix):

\(\large 2\tan\theta\) replaced \(\large u\). \(\large 2\sec^2\theta \;d\theta\) replaced \(\large du\).

OpenStudy (anonymous):

mkay i'm trying to get this all into this tiny space give me one moment to catch up

OpenStudy (anonymous):

I don't get why she gave us this tiny space to fit this huge problem in, alright continue from 1/2∫ dθ= 1/2 θ+C

zepdrix (zepdrix):

So we have to look back at the substitution we made, and try to solve it for \(\large \theta\). \[\large u=2\tan \theta \qquad \rightarrow \qquad \frac{u}{2}=\tan \theta \qquad \rightarrow \qquad \color{orangered}{\arctan\left(\frac{u}{2}\right)=\theta}\]There, we've successfully solved for theta. \[\large \frac{1}{2}\color{orangered}{\theta}+C \qquad = \qquad\frac{1}{2}\color{orangered}{\arctan\left(\frac{u}{2}\right)}+C\]

zepdrix (zepdrix):

Now we need to go from \(\large u\) to \(\large x\). Looking back at our substitution, \(\large \color{royalblue}{u=\sec x}\), \[\large \frac{1}{2}\arctan\left(\frac{\color{royalblue}{u}}{2}\right)+C \qquad = \qquad \frac{1}{2}\arctan\left(\frac{\color{royalblue}{\sec x}}{2}\right)+C\]

zepdrix (zepdrix):

If your teacher didn't give you much room, then she probably just wanted you to memorize the simple formula at this point :3 that would be my guess.

OpenStudy (anonymous):

I know we weren't taught this. Augh. It's a lot to take in. You have no idea how much help you are.

OpenStudy (anonymous):

If I could give more medals I would. thank you thank you thank you!

zepdrix (zepdrix):

heh np :3